# Information-theoretic minimax and submodular optimization algorithms for multivariate Markov chains

### v1

@@ 7.1.1 Curie-Weiss model
Let the Hamiltonian function be that of the Curie-Weiss model (see Chapter 13 of [22]) on $\mathcal{X}$ with interaction coefficients $\frac{1}{2^{|j-i|}}$ and external magnetic field $h=1$, that is, for $x=(x^{1},\ldots,x^{d})\in\mathcal{X}$,
@@ 7.1.2 Bernoulli-Laplace level model
We aim to generate a $d$-dimensional Markov chain from the Bernoulli-Laplace level model. We consider a $(d+1)$-dimensional Bernoulli–Laplace level model as described in Section 4.2 of [23]. Let
@@
be the state space, where $x^{i}$ can be interpreted as the number of “particles” of type $i$ out of the total number $N=d$. The stationary distribution of such Markov chain, $\pi$, is given by the multivariate hypergeometric distribution described in Lemma 4.18 of [23]. Concretely, we have
@@
Following the spectral decomposition for reversible Markov chains (see Section 2.1 of [23] for background), the transition matrix $P$ is written as:
@@
From Definition 4.15 of [23], in the Bernoulli-Laplace level model, we choose $s=1$ as the swap size parameter satisfying
@@
where we consider $\sum_{i=1}^{d+1}l_{i}>N$. From Theorem 4.19 of [23], the eigenvalues for the Bernoulli-Laplace level model are given by
@@
where $\mathbf{Q_{n}}$ are the multivariate Hahn polynomials for the hypergeometric distribution as defined in Proposition 2.3 of [23].
@@ 7.2 Numerical experiments of Algorithm 1
We then conduct experiments associated with the family of transition matrices including lazy Markov chain (see e.g. [24] for background). Precisely, we choose
@@ 7.3 Numerical experiments of Algorithm 2
For the Curie-Weiss model (Figure 6(a)), the objective value $f(\mathbf{S}_{i},\overline{\mathbf{w}}_{i})$ is not monotonically non-decreasing, as both the generalized distorted greedy algorithm (Algorithm 3 of [12]) and the projected subgradient algorithm (Algorithm 1) do not guarantee monotonicity. The final partition set is $\mathbf{S}_{l}=\mathbf{V}$, which means that the algorithm selects the ground set as the subset. After the final round of Algorithm 2, the final weight is $\overline{\mathbf{w}}_{l}=(0.70,0.00,0.00,0.00,0.30)$, which concentrates on the base transition matrix $P$ and the matrix with fastest mixing $P^{16}$.

### v2

@@ 7.1.1 Curie-Weiss model
Let the Hamiltonian function be that of the Curie-Weiss model (see Chapter 13 of [bovier2016metastability]) on $\mathcal{X}$ with interaction coefficients $\frac{1}{2^{|j-i|}}$ and external magnetic field $h=1$, that is, for $x=(x^{1},\ldots,x^{d})\in\mathcal{X}$,
@@ 7.1.2 Bernoulli-Laplace level model
We aim to generate a $d$-dimensional Markov chain from the Bernoulli-Laplace level model. We consider a $(d+1)$-dimensional Bernoulli–Laplace level model as described in Section 4.2 of [khare2009rates]. Let
@@
be the state space, where $x^{i}$ can be interpreted as the number of “particles” of type $i$ out of the total number $N=d$. The stationary distribution of such Markov chain, $\pi$, is given by the multivariate hypergeometric distribution described in Lemma 4.18 of [khare2009rates]. Concretely, we have
@@
Following the spectral decomposition for reversible Markov chains (see Section 2.1 of [khare2009rates] for background), the transition matrix $P$ is written as:
@@
From Definition 4.15 of [khare2009rates], in the Bernoulli-Laplace level model, we choose $s=1$ as the swap size parameter satisfying
@@
where we consider $\sum_{i=1}^{d+1}l_{i}>N$. From Theorem 4.19 of [khare2009rates], the eigenvalues for the Bernoulli-Laplace level model are given by
@@
where $\mathbf{Q_{n}}$ are the multivariate Hahn polynomials for the hypergeometric distribution as defined in Proposition 2.3 of [khare2009rates].
@@ 7.2 Numerical experiments of Algorithm 1
We then conduct experiments associated with the family of transition matrices including lazy Markov chain (see e.g. [shen2014lazy] for background). Precisely, we choose
@@ 7.3 Numerical experiments of Algorithm 2
For the Curie-Weiss model (Figure 6(a)), the objective value $f(\mathbf{S}_{i},\overline{\mathbf{w}}_{i})$ is not monotonically non-decreasing, as both the generalized distorted greedy algorithm (Algorithm 3 of [lai2025information]) and the projected subgradient algorithm (Algorithm 1) do not guarantee monotonicity. The final partition set is $\mathbf{S}_{l}=\mathbf{V}$, which means that the algorithm selects the ground set as the subset. After the final round of Algorithm 2, the final weight is $\overline{\mathbf{w}}_{l}=(0.70,0.00,0.00,0.00,0.30)$, which concentrates on the base transition matrix $P$ and the matrix with fastest mixing $P^{16}$.